# Positivity is undecidable in tensor products of free algebras

### 09 Dec 2023 v1

@@ 4.1. Group stage of construction
In the first stage of constructing $\mathcal{A}_{\mathcal{S}}(m)$, we build up a group $G_{\mathcal{S}}$ out of amalgamations, central products, and HNN extensions. We show that $G_{\mathcal{S}}$ has a finite generating set, and that the word problem with respect to this generating set is decidable in polynomial time. This allows us to embed $G_{\mathcal{S}}$ in a finitely presented group $H_{S}$, which we use as the basis for defining the algebras $\mathcal{A}_{\mathcal{S}}(m)$.
For the construction of $G_{\mathcal{S}}$, let $[A,B]:=ABA^{-1}B^{-1}$ denote the group commutator, and let $\mathbb{Z}_{n}$ be the additive cyclic group of order $n\in\mathbb{N}\cup\{+\infty\}$ (so in particular, $\mathbb{Z}_{\infty}=\mathbb{Z}$). To order the elements of $\mathbb{Z}_{n}$, we represent elements of $\mathbb{Z}_{n}$ by integers $-\lfloor\tfrac{n+2}{2}\rfloor<j<\lfloor\tfrac{n+1}{2}\rfloor$. For instance, for $\mathbb{Z}_{4}$ we use representatives $-2,-1,0,1$. This choice of representatives has the property that every congruence class is represented by an element that is minimal in absolute value, i.e. if $k^{\prime}$ is the representative of $[k]$, then $|k^{\prime}|\leq|k|$. Also, note that this set of representatives contains $-1$ for every $n\geq 2$. When $n=+\infty$, elements of $\mathbb{Z}_{n}=\mathbb{Z}$ represent themselves.
For each $m\in\mathbb{Z}$ and $i\in\mathbb{Z}_{h(m)+1}$, let $K_{\mathcal{S}}(m,i)$ be the group generated by indeterminates $J,x_{mi},z_{mi}$, subject to the relations:
@@
  • $[x_{mi},J]=[z_{mi},J]=1$, and
  • $[x_{mi},z_{mi}]=J^{a}$, where $a=1$ if $i{+}1\neq 0$ in $\mathbb{Z}_{h(m)+1}$, and $a=0$ otherwise.
@@
$K_{\mathcal{S}}(m,i)=\begin{cases}D_{4}&i+1\neq 0\text{ in }\mathbb{Z}_{h(m)+1}\\ \mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}&i+1=0\text{ in }\mathbb{Z}_{h(m)+1}\end{cases}.$
@@
$K_{\mathcal{S}}(m):=\sideset{{}_{\langle J\rangle}}{{}_{i\in\mathbb{Z}_{h(m)+1}}}{\prod}K_{\mathcal{S}}(m,i)$
be the central product of the groups $K_{\mathcal{S}}(m,i)$, $i\in\mathbb{Z}_{h(m)+1}$, over the central subgroup $\langle J\rangle=\mathbb{Z}_{2}$, and let
$K_{\mathcal{S}}:={}_{\langle J\rangle}\Asterisk_{m\in\mathbb{Z}}K_{\mathcal{S}}(m)$
be the amalgamated free product of the groups $K_{\mathcal{S}}(m)$, $m\in\mathbb{Z}$, over $\langle J\rangle=\mathbb{Z}_{2}$. In other words, $K_{\mathcal{S}}$ is generated by indeterminates
@@
    $[x_{mi},z_{mi}]=\begin{cases}J&i+1\neq 0\text{ in }\mathbb{Z}_{h(m)+1}\\ 1&i+1=0\text{ in }\mathbb{Z}_{h(m)+1}\end{cases},\text{ and }$
@@
    $[x_{mi},z_{mj}]=[x_{mi},x_{mj}]=[z_{mi},z_{mj}]=1.$
Note from relations (K1) that $J$ is central in $K_{\mathcal{S}}$. For each $m\in\mathbb{Z}$ and $i\in\mathbb{Z}_{h(m)+1}$, the elements $\{1,x_{mi},z_{mi},x_{mi}z_{mi}\}$ form a set of right coset representatives for $\langle J\rangle$. By the normal form theorem for central products, every element $w\in K_{\mathcal{S}}(m)$, $m\in\mathbb{Z}$ can be written uniquely as
@@
where $k\geq 0$, $i_{1}<i_{2}<\ldots<i_{k}$ are representatives of elements from $\mathbb{Z}_{h(m)+1}$ as described above, $c\in\mathbb{Z}_{2}$, and $(a_{j},b_{j})\in\{(0,1),(1,0),(1,1)\}\subset\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ for all $1\leq j\leq k$. The function $\eta_{m}(w)=J^{c}x_{m,i_{1}}^{a_{1}}z_{m,i_{1}}^{b_{1}}\cdots x_{m,i_{k}}^{a_{k}}z_{m,i_{k}}^{b_{k}}$ is a normal form for $K_{\mathcal{S}}(m)$. Note that we take the indices $i_{1},\ldots,i_{k}$ to be from the chosen set of representatives for $\mathbb{Z}_{h(m)+1}$, rather than elements of $\mathbb{Z}_{h(m)+1}$, so $\eta(w)$ is a word over
@@
$\max\{r\geq 0:\text{one of }x_{m,\pm r}^{\pm 1},z_{m,\pm r}^{\pm 1}\text{ appears in }w\text{ for some }m\in\mathbb{Z}\}\cup\{0\}.$
@@
$\mathcal{C}(m):=\big{\{}x_{m,i_{1}}^{a_{1}}z_{m,i_{1}}^{b_{1}}\cdots x_{m,i_{k}}^{a_{k}}z_{m,i_{k}}^{b_{k}}:$ $k\geq 0,-\lfloor\tfrac{h(m)+3}{2}\rfloor<i_{1}<\ldots<i_{k}<\lfloor\tfrac{h(m)+2}{2}\rfloor,$
@@
is a set of right coset representatives for $\langle J\rangle$ in $K_{\mathcal{S}}(m)$. By the normal form theorem for amalgamated free products, every element $w\in K_{\mathcal{S}}$ can be written uniquely as
@@
where $\ell\geq 0$, $c\in\mathbb{Z}_{2}$, and $w_{j}\in\mathcal{C}(m_{j})\setminus\{1\}$ for some integers $m_{1},\ldots,m_{\ell}$ such that $m_{j}\neq m_{j+1}$ for all $11\leq j\leq\ell-$ 1. The function $\eta(w)=J^{c}w_{1}\cdots w_{\ell}$ is a normal form for $K_{\mathcal{S}}$. This normal form can be computed efficiently:
@@
Note that if $-\lfloor\tfrac{h(m)+3}{2}\rfloor<i<\lfloor\tfrac{h(m)+2}{2}\rfloor$, then $i+1=0$ in $\mathbb{Z}_{h(m)+1}$ if and only if $i=-1$. Using relations (K1)-(K3), we can reorder $w$ so that it can be written as in Eq. 4.2 for some $c,a_{1},b_{1},\ldots,a_{k},b_{k}\in\mathbb{Z}$ (rather than $\mathbb{Z}_{2}$). Using relations (K0), we can then reduce $c,a_{1},b_{1},\ldots,a_{k},b_{k}$ mod $2$ to put $w$ in normal form $\eta_{m}(w)=J^{c}\overline{w}$, where $\overline{w}\in\mathcal{C}(m)$. Since a list of integers can be sorted in polynomial time, this process takes polynomial time in the bit-length of $w$. Since every generator in $\overline{w}$ occurs in $w$, and no generator appears in $\overline{w}$ more than once, the length, bit-length, and maximum index of $\overline{w}$ are at most the length, bit-length, and maximum index of $w$ respectively.
If $w$ is a word over $\widetilde{\mathcal{X}}_{K}\cup\widetilde{\mathcal{X}}_{K}^{-1}$, then we can write $w=J^{c}w_{1}\cdots w_{\ell}$, where $c\in\mathbb{Z}_{2}$, $w_{j}$ is a word over $\widetilde{\mathcal{X}}_{K}(m_{j})\cup\widetilde{\mathcal{X}}_{K}(m_{j})^{-1}$ for $1\leq i\leq\ell$, and $m_{1},\ldots,m_{\ell}$ is some sequence of integers such that $m_{j}\neq m_{j+1}$ for $11\leq j\leq\ell-$ 1. Using the procedure in the last paragraph, we can put $w_{j}$ in normal form $J^{c_{j}}\overline{w}_{j}$ for $K_{\mathcal{S}}(m_{j})$, where $\overline{w}_{j}\in\mathcal{C}(m_{j})$ and $c_{j}\in\mathbb{Z}_{2}$. The length of $\overline{w}_{j}$ is at most the length of $w_{j}$, so the length of $\overline{w}_{1}\cdots\overline{w}_{\ell}$ is at most the length of $w_{1}\cdots w_{\ell}$, and the same is true of bit-length and maximum index. Let $c^{\prime}=c+\sum_{j}c_{j}$. If $\overline{w}_{j}\neq 1$ for all $1\leq j\leq\ell$, then $\eta(w)=J^{c^{\prime}}\overline{w}_{1}\cdots\overline{w}_{\ell}$. Otherwise, if $\overline{w}_{j}=1$ for some $1\leq j\leq\ell$, then we can repeat this procedure with $w^{\prime}=J^{c^{\prime}}\overline{w}_{1}\cdots\overline{w}_{j-1}\overline{w}_{j+1}\cdots\overline{w}_{\ell}$. Since $\ell$ decreases in each iteration of this algorithm, the overall running time will be polynomial in the bit-length of $w$. Also, since the length of $w_{1}\cdots w_{\ell}$ does not increase in each iteration, the length of $\overline{w}_{1}\cdots\overline{w}_{\ell}$ in the final iteration is at most the length of the initial input $w$, and the same is true of bit-length and maximum index. ∎
@@
The above algorithm for computing $\eta$ assumes that the generators $x_{mi}$ and $z_{mi}$ in the input are given by specifying the representative of the index $i\in\mathbb{Z}_{h(m)+1}$. The algorithm does not work as stated without this assumption. For instance, on $w=x_{m,-3}\cdot x_{m,7}$ the algorithm will return $w$ back again, but if $h(m)=9$ then $x_{m,7}=x_{m,-3}$ in $K_{\mathcal{S}}$, and $\eta(w)=1$. If $i\in\mathbb{Z}$, then the representative for $i$ in $\mathbb{Z}_{h(m)+1}$ can be computed by running the Turing machine $M_{\mathcal{S}}$ on input $m$ for $12|i|-$ 1 steps. If $M_{\mathcal{S}}$ halts by this point then we know $h(m)$ and can replace $i$ by its representative if necessary, while if it does not halt then $h(m)>2|i|-1$ and hence $-\lfloor\frac{h(m)+3}{2}\rfloor<i<\lfloor\frac{h(m)+2}{2}\rfloor$. The running time of this procedure is polynomial in the bit-length of $m$ and the value of $i$, but not polynomial in the bit-length of $i$. We will use this procedure for computing the representative of $i$ later.
For the final step of constructing $G_{\mathcal{S}}$, we take iterated HNN extensions over several subgroups of $K_{\mathcal{S}}$:
$K_{\mathcal{S},x}$ $:=\langle J,x_{mi},m\in\mathbb{Z},i\in\mathbb{Z}_{h(m)+1}\rangle,$
$K_{\mathcal{S},z}$ $:=\langle J,z_{mi},m\in\mathbb{Z},i\in\mathbb{Z}_{h(m)+1}\rangle,\text{ and}$
$K_{\mathcal{S},0}$ $:=\langle J,x_{m,0},z_{m,0},m\in\mathbb{Z}\rangle.$
@@
    An element $w\in K_{\mathcal{S}}$ is in $K_{\mathcal{S},x}$ if and only if its normal form $\eta(w)$ is a word over $\{J,x_{mi},\ m\in\mathbb{Z},-\lfloor\tfrac{h(m)+3}{2}\rfloor<i<\lfloor\tfrac{h(m)+2}{2}\rfloor\}$. In addition, there is an automorphism $\phi_{S}:K_{\mathcal{S},x}\rightarrow K_{\mathcal{S},x}$ sending $J\mapsto J$ and $x_{mi}\mapsto x_{m,i+1}$ for all $m\in\mathbb{Z},i\in\mathbb{Z}_{h(m)+1}$.
@@
    An element $w\in K_{\mathcal{S}}$ is in $K_{\mathcal{S},z}$ if and only if its normal form $\eta(w)$ is a word over $\{J,z_{mi},\ m\in\mathbb{Z},-\lfloor\tfrac{h(m)+3}{2}\rfloor<i<\lfloor\tfrac{h(m)+2}{2}\rfloor\}$. In addition, there is an automorphism $\phi_{T}:K_{\mathcal{S},z}\rightarrow K_{\mathcal{S},z}$ sending $J\mapsto J$ and   $z_{mi}\mapsto z_{m,i+1}$ for all $m\in\mathbb{Z},i\in\mathbb{Z}_{h(m)+1}$.
@@
    An element $w\in K_{\mathcal{S}}$ is in $K_{\mathcal{S},0}$ if and only if its normal form $\eta(w)$ is a word over $\{J,x_{m,0},z_{m,0},m\in\mathbb{Z},\}$. In addition, there is an automorphism $\phi_{W}:K_{\mathcal{S},0}\rightarrow K_{\mathcal{S},0}$ sending $J\mapsto J$, $x_{m,0}\mapsto x_{m+1,0}$, and $z_{m,0}\mapsto z_{m+1,0}$ for all $m\in\mathbb{Z}$.
@@
For part (a), observe that the subgroup $K_{\mathcal{S},x}(m,i):=\langle J,x_{mi}\rangle\subseteq K_{\mathcal{S}}(m,i)$ is isomorphic to $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ for all $m\in\mathbb{Z},i\in\mathbb{Z}_{h(m)+1}$. For $m\in\mathbb{Z}$, let $K_{\mathcal{S},x}(m):=\langle J,x_{mi}:i\in\mathbb{Z}_{h(m)+1}\rangle\subseteq K_{\mathcal{S}}(m)$, and consider the natural surjective homomorphism
$\Phi_{m}:\sideset{{}_{\langle J\rangle}}{{}_{i\in\mathbb{Z}_{h(m)+1}}}{\prod}K_{\mathcal{S},x}(m,i)\rightarrow K_{\mathcal{S},x}(m).$
By the normal form theorem for central products, every element $w$ in the domain of $\Phi_{m}$ can be written uniquely as $w=J^{c}x_{m,i_{1}}x_{m,i_{2}}\cdots x_{m,i_{k}}$, where $k\geq 0$, $-\lfloor\tfrac{h(m)+3}{2}\rfloor<i_{1}<\cdots<i_{k}<\lfloor\tfrac{h(m)+2}{2}\rfloor$, and $c\in\mathbb{Z}_{2}$. Comparing with the normal form $\eta_{m}$ for $K_{\mathcal{S}}(m)$, we see that $\Phi_{m}$ sends normal forms to normal forms. It follows that $\Phi_{m}$ is an isomorphism, and that if $w\in K_{\mathcal{S},x}(m)$, then the normal form $\eta_{m}(w)$ in $K_{\mathcal{S}}(m)$ is a word containing only $J$ and $x_{mi}$’s.
@@
$K_{\mathcal{S},x}={}_{\langle J\rangle}\Asterisk_{m\in\mathbb{Z}}K_{\mathcal{S},x}(m),$
and the normal form $\eta(w)$ of an element $w\in K_{\mathcal{S},x}$ is a word over $\{J,x_{mi}:m\in\mathbb{Z},-\lfloor\tfrac{h(m)+3}{2}\rfloor<i<\lfloor\tfrac{h(m)+2}{2}\rfloor\}$.

### 18 Mar 2026 v2

@@ 4.1. Group stage of construction
In the first stage of constructing $\mathcal{A}_{\mathcal{L}}(m)$, we build up a group $G_{\mathcal{L}}$ out of amalgamations, central products, and HNN extensions. We show that $G_{\mathcal{L}}$ has a finite generating set, and that the word problem with respect to this generating set is decidable in polynomial time. This allows us to embed $G_{\mathcal{L}}$ in a finitely presented group $H_{S}$, which we use as the basis for defining the algebras $\mathcal{A}_{\mathcal{L}}(m)$.
For the construction of $G_{\mathcal{L}}$, let $A,B]:=ABA^{-1}B^{-1}$ denote the group commutator, and let $\mathbb{Z}_{n}$ be the additive cyclic group of order $n\in\mathbb{N}\cup\{+\infty\}$ (so in particular, $\mathbb{Z}_{\infty}=\mathbb{Z}$). To order the elements of $\mathbb{Z}_{n}$, we represent elements of $\mathbb{Z}_{n}$ by integers $\lfloor\tfrac{n+2}{2}\rfloor<j<\lfloor\tfrac{n+1}{2}\rfloor$. For instance, for $\mathbb{Z}_{4}$ we use representatives 2,1,0,1-2,-1,0,1. This choice of representatives has the property that every congruence class is represented by an element that is minimal in absolute value, i.e. if $k^{\prime}$ is the representative of $[k]$, then $|k^{\prime}|\leq|k|$. Also, note that this set of representatives contains 1-1 for every $n\geq 2$. When $n=+\infty$, elements of $\mathbb{Z}_{n}=\mathbb{Z}$ represent themselves.
For each $m\in\mathbb{Z}$ and $i\in\mathbb{Z}_{h(m)+1}$, let $K_{\mathcal{L}}(m,i)$ be the group generated by indeterminates $J,x_{mi},z_{mi}$, subject to the relations:
@@
  • $x_{mi},J]=[z_{mi},J]=1$, and
  • $x_{mi},z_{mi}]=J^{a}$, where $a=1$ if $i{+}1\neq 0$ in $\mathbb{Z}_{h(m)+1}$, and $a=0$ otherwise.
@@
$K_{\mathcal{L}}(m,i)=\begin{cases}D_{4}&i+1\neq 0\text{ in }\mathbb{Z}_{h(m)+1}\\ \mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}&i+1=0\text{ in }\mathbb{Z}_{h(m)+1}\end{cases}.$
@@
$K_{\mathcal{L}}(m):=\sideset{{}_{\langle J\rangle}}{{}_{i\in\mathbb{Z}_{h(m)+1}}}{\prod}K_{\mathcal{L}}(m,i)$
be the central product of the groups $K_{\mathcal{L}}(m,i)$, $i\in\mathbb{Z}_{h(m)+1}$, over the central subgroup $\langle J\rangle=\mathbb{Z}_{2}$, and let
$K_{\mathcal{L}}:={}_{\langle J\rangle}\Asterisk_{m\in\mathbb{Z}}K_{\mathcal{L}}(m)$
be the amalgamated free product of the groups $K_{\mathcal{L}}(m)$, $m\in\mathbb{Z}$, over $\langle J\rangle=\mathbb{Z}_{2}$. In other words, $K_{\mathcal{L}}$ is generated by indeterminates
@@
    $x_{mi},z_{mi}]=\begin{cases}J&i+1\neq 0\text{ in }\mathbb{Z}_{h(m)+1}\\ 1&i+1=0\text{ in }\mathbb{Z}_{h(m)+1}\end{cases},\text{ and }$
@@
    $x_{mi},z_{mj}]=[x_{mi},x_{mj}]=[z_{mi},z_{mj}]=1.$
Note from relations (K1) that $J$ is central in $K_{\mathcal{L}}$. For each $m\in\mathbb{Z}$ and $i\in\mathbb{Z}_{h(m)+1}$, the elements $\{1,x_{mi},z_{mi},x_{mi}z_{mi}\}$ form a set of right coset representatives for $\langle J\rangle$. By the normal form theorem for central products, every element $w\in K_{\mathcal{L}}(m)$, $m\in\mathbb{Z}$ can be written uniquely as
@@
where $k\geq 0$, $i_{1}<i_{2}<\ldots<i_{k}$ are representatives of elements from $\mathbb{Z}_{h(m)+1}$ as described above, $c\in\mathbb{Z}_{2}$, and $(a_{j},b_{j})\in\{(0,1),(1,0),(1,1)\}\subset\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ for all $1\leq j\leq k$. The function $\eta_{m}(w)=J^{c}x_{m,i_{1}}^{a_{1}}z_{m,i_{1}}^{b_{1}}\cdots x_{m,i_{k}}^{a_{k}}z_{m,i_{k}}^{b_{k}}$ is a normal form for $K_{\mathcal{L}}(m)$. Note that we take the indices $i_{1},\ldots,i_{k}$ to be from the chosen set of representatives for $\mathbb{Z}_{h(m)+1}$, rather than elements of $\mathbb{Z}_{h(m)+1}$, so $\eta(w)$ is a word over
@@
[…]
@@
$\mathcal{C}(m):=\big\{x_{m,i_{1}}^{a_{1}}z_{m,i_{1}}^{b_{1}}\cdots x_{m,i_{k}}^{a_{k}}z_{m,i_{k}}^{b_{k}}:\$$k\geq 0,-\lfloor\tfrac{h(m)+3}{2}\rfloor<i_{1}<\ldots<i_{k}<\lfloor\tfrac{h(m)+2}{2}\rfloor,$
@@
is a set of right coset representatives for $\langle J\rangle$ in $K_{\mathcal{L}}(m)$. By the normal form theorem for amalgamated free products, every element $w\in K_{\mathcal{L}}$ can be written uniquely as
@@
where $\ell\geq 0$, $c\in\mathbb{Z}_{2}$, and $w_{j}\in\mathcal{C}(m_{j})\setminus\{1\}$ for some integers $m_{1},\ldots,m_{\ell}$ such that $m_{j}\neq m_{j+1}$ for all $1\leq j\leq\ell-1$. The function $\eta(w)=J^{c}w_{1}\cdots w_{\ell}$ is a normal form for $K_{\mathcal{L}}$. This normal form can be computed efficiently:
@@
Note that if $\lfloor\tfrac{h(m)+3}{2}\rfloor<i<\lfloor\tfrac{h(m)+2}{2}\rfloor$, then $i+1=0$ in $\mathbb{Z}_{h(m)+1}$ if and only if i=1i=-1. Using relations (K1)-(K3), we can reorder $w$ so that it can be written as in Eq. 4.2 for some $c,a_{1},b_{1},\ldots,a_{k},b_{k}\in\mathbb{Z}$ (rather than $\mathbb{Z}_{2}$). Using relations (K0), we can then reduce $c,a_{1},b_{1},\ldots,a_{k},b_{k}$ mod $2$ to put $w$ in normal form $\eta_{m}(w)=J^{c}\overline{w}$, where $\overline{w}\in\mathcal{C}(m)$. Since a list of integers can be sorted in polynomial time, this process takes polynomial time in the bit-length of $w$. Since every generator in $\overline{w}$ occurs in $w$, and no generator appears in $\overline{w}$ more than once, the length, bit-length, and maximum index of $\overline{w}$ are at most the length, bit-length, and maximum index of $w$ respectively.
If $w$ is a word over $\widetilde{\mathcal{X}}_{K}\cup\widetilde{\mathcal{X}}_{K}^{-1}$, then we can write $w=J^{c}w_{1}\cdots w_{\ell}$, where $c\in\mathbb{Z}_{2}$, $w_{j}$ is a word over $\widetilde{\mathcal{X}}_{K}(m_{j})\cup\widetilde{\mathcal{X}}_{K}(m_{j})^{-1}$ for $1\leq i\leq\ell$, and $m_{1},\ldots,m_{\ell}$ is some sequence of integers such that $m_{j}\neq m_{j+1}$ for $1\leq j\leq\ell-1$. Using the procedure in the last paragraph, we can put $w_{j}$ in normal form $J^{c_{j}}\overline{w}_{j}$ for $K_{\mathcal{L}}(m_{j})$, where $\overline{w}_{j}\in\mathcal{C}(m_{j})$ and $c_{j}\in\mathbb{Z}_{2}$. The length of $\overline{w}_{j}$ is at most the length of $w_{j}$, so the length of $\overline{w}_{1}\cdots\overline{w}_{\ell}$ is at most the length of $w_{1}\cdots w_{\ell}$, and the same is true of bit-length and maximum index. Let $c^{\prime}=c+\sum_{j}c_{j}$. If $\overline{w}_{j}\neq 1$ for all $1\leq j\leq\ell$, then $\eta(w)=J^{c^{\prime}}\overline{w}_{1}\cdots\overline{w}_{\ell}$. Otherwise, if $\overline{w}_{j}=1$ for some $1\leq j\leq\ell$, then we can repeat this procedure with $w^{\prime}=J^{c^{\prime}}\overline{w}_{1}\cdots\overline{w}_{j-1}\overline{w}_{j+1}\cdots\overline{w}_{\ell}$. Since $\ell$ decreases in each iteration of this algorithm, the overall running time will be polynomial in the bit-length of $w$. Also, since the length of $w_{1}\cdots w_{\ell}$ does not increase in each iteration, the length of $\overline{w}_{1}\cdots\overline{w}_{\ell}$ in the final iteration is at most the length of the initial input $w$, and the same is true of bit-length and maximum index. ∎
@@
The above algorithm for computing $\eta$ assumes that the generators $x_{mi}$ and $z_{mi}$ in the input are given by specifying the representative of the index $i\in\mathbb{Z}_{h(m)+1}$. The algorithm does not work as stated without this assumption. For instance, on $w=x_{m,-3}\cdot x_{m,7}$ the algorithm will return $w$ back again, but if $h(m)=9$ then $x_{m,7}=x_{m,-3}$ in $K_{\mathcal{L}}$, and $\eta(w)=1$. If $i\in\mathbb{Z}$, then the representative for $i$ in $\mathbb{Z}_{h(m)+1}$ can be computed by running the Turing machine $M_{\mathcal{L}}$ on input $m$ for 2|i|12|i|-1 steps. If $M_{\mathcal{L}}$ halts by this point then we know $h(m)$ and can replace $i$ by its representative if necessary, while if it does not halt then h(m)>2|i|1h(m)>2|i|-1 and hence $\lfloor\frac{h(m)+3}{2}\rfloor<i<\lfloor\frac{h(m)+2}{2}\rfloor$. The running time of this procedure is polynomial in the bit-length of $m$ and the value of $i$, but not polynomial in the bit-length of $i$. We will use this procedure for computing the representative of $i$ later.
For the final step of constructing $G_{\mathcal{L}}$, we take iterated HNN extensions over several subgroups of $K_{\mathcal{L}}$:
$K_{\mathcal{L},x}$ $:=\langle J,x_{mi},m\in\mathbb{Z},i\in\mathbb{Z}_{h(m)+1}\rangle,$
$K_{\mathcal{L},z}$ $:=\langle J,z_{mi},m\in\mathbb{Z},i\in\mathbb{Z}_{h(m)+1}\rangle,\text{ and}$
$K_{\mathcal{L},0}$ $:=\langle J,x_{m,0},z_{m,0},m\in\mathbb{Z}\rangle.$
@@
    An element $w\in K_{\mathcal{L}}$ is in $K_{\mathcal{L},x}$ if and only if its normal form $\eta(w)$ is a word over $\{J,x_{mi},\ m\in\mathbb{Z},-\lfloor\tfrac{h(m)+3}{2}\rfloor<i<\lfloor\tfrac{h(m)+2}{2}\rfloor\}$. In addition, there is an automorphism $\phi_{S}:K_{\mathcal{L},x}\rightarrow K_{\mathcal{L},x}$ sending $J\mapsto J$ and $x_{mi}\mapsto x_{m,i+1}$ for all $m\in\mathbb{Z},i\in\mathbb{Z}_{h(m)+1}$.
@@
    An element $w\in K_{\mathcal{L}}$ is in $K_{\mathcal{L},z}$ if and only if its normal form $\eta(w)$ is a word over $\{J,z_{mi},\ m\in\mathbb{Z},-\lfloor\tfrac{h(m)+3}{2}\rfloor<i<\lfloor\tfrac{h(m)+2}{2}\rfloor\}$. In addition, there is an automorphism $\phi_{T}:K_{\mathcal{L},z}\rightarrow K_{\mathcal{L},z}$ sending $J\mapsto J$ and   $z_{mi}\mapsto z_{m,i+1}$ for all $m\in\mathbb{Z},i\in\mathbb{Z}_{h(m)+1}$.
@@
    An element $w\in K_{\mathcal{L}}$ is in $K_{\mathcal{L},0}$ if and only if its normal form $\eta(w)$ is a word over $\{J,x_{m,0},z_{m,0},m\in\mathbb{Z}\}$. In addition, there is an automorphism $\phi_{W}:K_{\mathcal{L},0}\rightarrow K_{\mathcal{L},0}$ sending $J\mapsto J$, $x_{m,0}\mapsto x_{m+1,0}$, and $z_{m,0}\mapsto z_{m+1,0}$ for all $m\in\mathbb{Z}$.
@@
For part (a), observe that the subgroup $K_{\mathcal{L},x}(m,i):=\langle J,x_{mi}\rangle\subseteq K_{\mathcal{L}}(m,i)$ is isomorphic to $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ for all $m\in\mathbb{Z},i\in\mathbb{Z}_{h(m)+1}$. For $m\in\mathbb{Z}$, let $K_{\mathcal{L},x}(m):=\langle J,x_{mi}:i\in\mathbb{Z}_{h(m)+1}\rangle\subseteq K_{\mathcal{L}}(m)$, and consider the natural surjective homomorphism
$\Phi_{m}:\sideset{{}_{\langle J\rangle}}{{}_{i\in\mathbb{Z}_{h(m)+1}}}{\prod}K_{\mathcal{L},x}(m,i)\rightarrow K_{\mathcal{L},x}(m).$
By the normal form theorem for central products, every element $w$ in the domain of $\Phi_{m}$ can be written uniquely as $w=J^{c}x_{m,i_{1}}x_{m,i_{2}}\cdots x_{m,i_{k}}$, where $k\geq 0$, $\lfloor\tfrac{h(m)+3}{2}\rfloor<i_{1}<\cdots<i_{k}<\lfloor\tfrac{h(m)+2}{2}\rfloor$, and $c\in\mathbb{Z}_{2}$. Comparing with the normal form $\eta_{m}$ for $K_{\mathcal{L}}(m)$, we see that $\Phi_{m}$ sends normal forms to normal forms. It follows that $\Phi_{m}$ is an isomorphism, and that if $w\in K_{\mathcal{L},x}(m)$, then the normal form $\eta_{m}(w)$ in $K_{\mathcal{L}}(m)$ is a word containing only $J$ and $x_{mi}$’s.
@@
$K_{\mathcal{L},x}={}_{\langle J\rangle}\Asterisk_{m\in\mathbb{Z}}K_{\mathcal{L},x}(m),$
and the normal form $\eta(w)$ of an element $w\in K_{\mathcal{L},x}$ is a word over $\{J,x_{mi}:m\in\mathbb{Z},-\lfloor\tfrac{h(m)+3}{2}\rfloor<i<\lfloor\tfrac{h(m)+2}{2}\rfloor\}$.